OpenStudy (anonymous):

Give the equation of the line that passes through the points (-3,-3) and (-4,2)

pooja195 (pooja195):

First find the slope \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\] also what form do they want it in?

OpenStudy (anonymous):

y =

pooja195 (pooja195):

y=mx+b?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its actually ) -3,-4) and (-4,2)